A 6730 - kg truck traveling at 6.1 m/s east collides with a 2470 - kg car moving at 23.66 m/s in a direction 35° south of west. After the collision, the two vehicles remain tangled together. With what velocity (magnitude only) does the wreckage begin to move?


1
Expert's answer
2021-03-22T08:36:18-0400

Assuming the collision as perfectly elastic,

Kinetic energy before collision equals to the kinetic energy after collision.

Let,

m1 = mass of truck= 6730 kg

m2 = mass of car 2470 kg

V1 = speed of truck=6.1 m/s

V2 = speed of car 23.33 m/s

V = final velocity of combined wreckage after collision

By conservation of energy,

"\frac{m_1×V^2_1}{2}+\frac{m_2×V^2_2}{2}=(m_1+m_2). \frac{V^2}{2}"

V=86.674 m/s

Therefore wreckage moves with speed 86.674 m/s